OpenStudy (anonymous):

If f"(x)-f'(x)-2f(x)=0, f'(0)=-2, and f(0)=2, what's f(1)? (Answer: 2e^-1)

OpenStudy (anonymous):

y'' - y' -2 y=0 Assumes linear equation hence f(x)=A e^x f(x)= A e^cx f(0)= A e^c0 =A*1=2 A=2 f'(x)=cA e^cx f'(0)= cA e^(c0)=-2 cA=-2 A=2 hence c=-1

OpenStudy (anonymous):

now with what you know plug in 1 for f(x)

OpenStudy (loser66):

wonder how can you not know how to solve it.???? characteristic equation r^2 - r -2 = 0 r =2 and r =-1 general solution is \[f(x)= C_1e^{-x}+C_2e^{2x}\] \[f(0) = C_1 +C_2 = 2~~~ (*)\] \[f'(x) = -C_1e^{-x}+2C_2e^{2x}\] \[f'(0) = -C_1 +2C_2=-2~~~(**)\] (*) and (**) ---> \(C_1 = 2\) and \(C_2 =0\) So, the solution is \(f(x) = 2e^{-x}\) replace x = -1 \(f(-1) = 2e\) I am sorry for getting the solution which is not as yours but I don't know what my mistake is

OpenStudy (anonymous):

my solution is valid and would have gotten your answer. how can you be so ungrateful as to ignore my response.